OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, -8) and a directrix at y = 8.

OpenStudy (anonymous):

@BulletWithButterflyWings

OpenStudy (anonymous):

@lncognlto do you get it? or are you able to explain it to me?

OpenStudy (lncognlto):

I'm sorry, Chandler, I don't know about focus and directx off the top of my head. I can look it up, but it might take a little while... Do you want me to?

OpenStudy (anonymous):

Eh I'm sure there is one person that does. Don't worry about it. Thanks though! @lncognlto

OpenStudy (anonymous):

@Shikanu

OpenStudy (lncognlto):

Cool cool :)

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@nincompoop

OpenStudy (campbell_st):

the general form is \[x^2 = -4ay\] since the directrix is above the focus well the distance between the focus and directrix is 2a so 16 = 2a, so a = 8 substitute it into the equation. and you get the parabola

OpenStudy (anonymous):

Thank you!! @campbell_st

OpenStudy (anonymous):

so wait I will get x^2 = -4(8)y ?

OpenStudy (anonymous):

giving me x^2 = -32y?

OpenStudy (anonymous):

Now what?... @campbell_st

OpenStudy (campbell_st):

thats the equation or it can be written as y = -1/32 x^2